# Convex Geometry of Building Sets

### v3

@@ 4. Restrictions of building sets and nested set complexes
4.5 (1): Because $\hookrightarrow$ is a meet-semilattice order embedding, we can identify $\leq_{\mathcal{K}}$ with $\leq_{\mathcal{L}}$, and $\wedge^{\mathcal{K}}$ with $\wedge^{\mathcal{L}}$, which we will denote by $\leq$ and $\wedge$, respectively, throughout the proof. Let $\sigma$ be the building set closure operator associated to $\mathcal{K}$. We begin by proving that $\mathbb{B}(\mathcal{L})=\{\sigma(B)\cap\mathcal{L}^{+}:B\in\mathbb{B}(\mathcal{L})\}.$ Let $\mathbb{X}=\{B\in\mathbb{B}(\mathcal{L}):\sigma(B)\cap\mathcal{L}^{+}=B\}.$ We would like to show that $\mathbb{X}=\mathbb{B}(\mathcal{L})$. Suppose for contradiction that there exists some $A\in\mathbb{B}(\mathcal{L})\setminus\mathbb{X}$. We know that $\mathbb{X}$ is non empty as $\mathcal{K}^{+}\cap\mathcal{L}^{+}=\mathcal{L}^{+}\in\mathbb{X}$, hence there exists a containment minimal element $B\in\mathbb{X}$ such that $A\subsetneq B$. By the supersolvability of $\mathbb{B}(\mathcal{L})$ applied to the pair of building sets $A$ and $B$, we know that for any linear extension $<$ of $\mathcal{L}$ and $x=\text{min}_{<}(B\setminus A)$ that $B\setminus\{x\}\in\mathbb{B}(\mathcal{L})$. By our minimality assumption on $B$, we have $\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}\neq B\setminus\{x\}$. So there exists some element $y\in\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}$ and $y\notin B\setminus\{x\}$. By condition (2) of Definition 3.2, we have that $B\setminus\{x\}\subsetneq\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}\subseteq\sigma(B)\cap\mathcal{L}^{+}=B$. Thus $\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}=B$ and $y=x$.
@@
The argument above establishes that $\mathbb{B}(\mathcal{L})\subseteq\{B\cap\mathcal{L}^{+}:B\in\mathbb{B}(\mathcal{K}),I(\mathcal{L})\subseteq B\}$, thus for completing the proof of (1) from Theorem 4.5, we will demonstrate the opposite inclusion. Let $\mathbb{Y}=\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\,,\,B\cap\mathcal{L}^{+}\in\mathbb{B}(\mathcal{L})\}.$ We will show that $\mathbb{Y}=\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\}$. Suppose for contradiction that $\mathbb{Y}\neq\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\}$ and let $A\in\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\}\setminus\mathbb{Y}$. We know that $\mathcal{K}^{+}\in\mathbb{Y}$, so $\mathbb{Y}$ is non empty, hence there exists a containment minimal building set $B\in\mathbb{Y}$ such that $A\subsetneq B$. By the supersolvability of $\mathbb{B}(\mathcal{K})$ applied to the pair of building sets $A$ and $B$, we know that for any linear extension $<$ of $\mathcal{K}$ and $x=\text{min}_{<}(B\setminus A)$, we have $B\setminus\{x\}\in\mathbb{B}(\mathcal{K})$. By our minimality assumption on $B$, we have that $(B\setminus\{x\})\cap\mathcal{L}^{+}\notin\mathbb{B}(\mathcal{L})$, i.e. $x$ is not an extreme point of $B$ in $\mathcal{L}$. We know that $I(\mathcal{L})\subseteq A$, hence $x\notin I(\mathcal{L})$ and $I(\mathcal{L})\subseteq(B\setminus\{x\})\cap\mathcal{L}^{+}$. By Proposition 3.15, there must exist some $y,z\in(B\setminus\{x\})\cap\mathcal{L}^{+}$ such that $y\wedge z\neq\hat{0}$ and $y\vee^{\mathcal{L}}z=x$. We know that $B\setminus\{x\}\in\mathbb{B}(\mathcal{K})$. Take $\{w_{1},\ldots,w_{k}\}=\text{max}\,(B\setminus\{x\})_{\leq x}$, and recall that for each $1\leq i<j\leq k$, we have that $[\hat{0},w_{i}]\cap[\hat{0},w_{j}]=\hat{0}$. Note that since $x\notin B\setminus\{x\}$, we must have $k\geq 2$. Also, if $u\in(B\setminus\{x\})_{\leq x}$ and $u\wedge w_{i}\neq\hat{0}$, then $u\leq w_{i}$. As $y,z\in B\setminus\{x\}$ and $\hat{0}\neq y\wedge z$, we know that $y\vee^{\mathcal{K}}z\in B\setminus\{x\}$. We may assume without loss of generality that $y\vee^{\mathcal{K}}z\leq w_{1}$. Let $\{x_{1},\ldots,x_{t}\}=\text{max}\,I(\mathcal{L})_{\leq x}\subseteq B\setminus\{x\}$. Suppose that for some $s$ with $1\leq s\leq t$, we have $x_{s}\wedge y=x_{s}\wedge z=\hat{0}$. Because $I(\mathcal{L})\in\mathbb{B}(\mathcal{L})$, we know that for $1\leq i\leq t$ there exist $y_{i},z_{i}\leq x_{i}$ such that $\vee_{1\leq i\leq t}^{\mathcal{L}}y_{i}=y$ and $\vee_{1\leq i\leq t}^{\mathcal{L}}z_{i}=z$. By assumption, $y_{s}=z_{s}=\hat{0}$, so $y\vee^{\mathcal{L}}z\leq\vee_{i\neq s}^{\mathcal{L}}x_{i}<\vee_{1\leq i\leq t}^{\mathcal{L}}x_{i}=x$, a contradiction. Thus for each $i$, the intersection $x_{i}\wedge y\neq\hat{0}$ or $x_{i}\wedge z\neq\hat{0}$, hence $x_{i}\wedge w_{1}\neq\hat{0}$. As $x_{i}\in(B\setminus\{x\})_{\leq x}$, we must have $x_{i}\leq w_{1}$, therefore $I(\mathcal{L})_{\leq x}\subseteq[\hat{0},w_{1}]$. Because $\mathcal{L}\hookrightarrow\mathcal{K}$ is consistent, $I(\mathcal{K})_{\leq x}\subseteq I(\mathcal{L})_{\leq x}$. We conclude that $I(\mathcal{K})_{\leq x}\subseteq[\hat{0},w_{1}]$, implying $k=1$, a contradiction.
@@
(2): Suppose for contradiction that $B\in\mathbb{B}(\mathcal{K})$, and $I(\mathcal{L})\subseteq B$, but $\mathcal{N}_{\mathcal{L}}(B\cap\mathcal{L}^{+})\nsubseteq\mathcal{N}_{\mathcal{K}}(B).$ Let $N\in\mathcal{N}_{\mathcal{L}}(B\cap\mathcal{L}^{+})$ such that $N\notin\mathcal{N}_{\mathcal{K}}(B)$. There exists some $x_{1},\ldots,x_{k}\in N$ which are incomparable, such that $y=\bigvee_{1\leq i\leq k}^{\mathcal{K}}x_{i}\in B$ but $x=\bigvee_{1\leq i\leq k}^{\mathcal{L}}x_{i}\in\mathcal{L}^{+}\setminus B$. Let $\{y_{1},\ldots,y_{l}\}=\text{max}\,B_{\leq x}$. Then $l\geq 2$ and the intervals $[\hat{0},y_{i}]$ intersect at $\hat{0}$. By Lemma 2.4, $y\leq x$, thus there exists exactly one $i$ such that $y\leq y_{i}$. But then by Lemma 2.10 applied to $x_{1},\ldots,x_{k}$, we know that for any $x^{\prime}\in(B\cap\mathcal{L}^{+})_{\leq x}$, there exists some $x_{t}$ such that $x^{\prime}\leq x_{t}$, hence $x^{\prime}\leq y\leq y_{i}$. It follows that for any $y_{j}$ with $j\neq i$, we have that $(B\cap\mathcal{L}^{+})_{\leq y_{j}}=\emptyset$. On the other hand, there must exist some $y^{\prime}\in I(\mathcal{K})_{\leq y_{j}}$. Clearly, $I(\mathcal{K})_{\leq y_{j}}\subseteq I(\mathcal{K})_{\leq x}$, and by our assumption that $\mathcal{L}\hookrightarrow\mathcal{K}$ is consistent, we must have that $y^{\prime}\in I(\mathcal{L})_{\leq x}\subseteq(B\cap\mathcal{L}^{+})_{\leq y_{i}}$, a contradiction.
@@ 5. Acknowledgements
The authors were supported by NSF Grant (DMS-2246967) and Simons Collaboration Gift  #854037. We thank Nate Bottman, Federico Castillo, Graham Denham, Alexei Oblomkov, Arnau Padrol, Vincent Pilaud, Dasha Poliakova, and Raman Sanyal for helpful discussions about building sets, and we thank Bryan Gillespie for carefully explaining his work during the 2017 MSRI Semester on “Geometric and Topological Combinatorics”. Some of our discussions took place at the Simons Center for Geometry and Physics Workshop on “Combinatorics and Geometry of Convex Polyhedra” in March 2023. S.B. thanks the MSRI (now SLMath) and the Simons Center for their hospitality and excellent working conditions. We thank Christin Bibby, Graham Denham, Matt Larson, Arnau Pardol, Vincent Pilaud, Vic Reiner, and Raman Sanyal for their helpful feedback on an earlier version of this article.

### v4

@@ 4. Restrictions of building sets and nested set complexes
4.5 (1): Because $\hookrightarrow$ is a meet-semilattice order embedding, we can identify $\leq_{\mathcal{K}}$ with $\leq_{\mathcal{L}}$, and $\wedge^{\mathcal{K}}$ with $\wedge^{\mathcal{L}}$, which we will denote by $\leq$ and $\wedge$, respectively, throughout the proof. Let $\sigma$ be the building set closure operator associated to $\mathcal{K}$. We begin by proving that $\mathbb{B}(\mathcal{L})=\{\sigma(B)\cap\mathcal{L}^{+}:B\in\mathbb{B}(\mathcal{L})\}.$ Let $\mathbb{X}=\{B\in\mathbb{B}(\mathcal{L}):\sigma(B)\cap\mathcal{L}^{+}=B\}.$ We would like to show that $\mathbb{X}=\mathbb{B}(\mathcal{L})$. Suppose for contradiction that there exists some $A\in\mathbb{B}(\mathcal{L})\setminus\mathbb{X}$. We know that $\mathbb{X}$ is non empty as $\mathcal{K}^{+}\cap\mathcal{L}^{+}=\mathcal{L}^{+}\in\mathbb{X}$, hence there exists a containment minimal element $B\in\mathbb{X}$ such that $A\subsetneq B$. By the supersolvability of $\mathbb{B}(\mathcal{L})$ applied to the pair of building sets $A$ and $B$, we know that for any linear extension << of $\mathcal{L}$ and $x=\text{min}_{<}(B\setminus A)$ that $B\setminus\{x\}\in\mathbb{B}(\mathcal{L})$. By our minimality assumption on $B$, we have $\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}\neq B\setminus\{x\}$. So there exists some element $y\in\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}$ and $y\notin B\setminus\{x\}$. By condition (2) of Definition 3.2, we have that $B\setminus\{x\}\subsetneq\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}\subseteq\sigma(B)\cap\mathcal{L}^{+}=B$. Thus $\sigma(B\setminus\{x\})\cap\mathcal{L}^{+}=B$ and $y=x$.
@@
The argument above establishes that $\mathbb{B}(\mathcal{L})\subseteq\{B\cap\mathcal{L}^{+}:B\in\mathbb{B}(\mathcal{K}),I(\mathcal{L})\subseteq B\}$, thus to complete the proof of (1) from Theorem 4.5, we will demonstrate the opposite inclusion. Let $\mathbb{Y}=\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\,,\,B\cap\mathcal{L}^{+}\in\mathbb{B}(\mathcal{L})\}.$ We will show that $\mathbb{Y}=\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\}$. Suppose for contradiction that $\mathbb{Y}\neq\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\}$ and let $A\in\{B\in\mathbb{B}(\mathcal{K}):I(\mathcal{L})\subseteq B\}\setminus\mathbb{Y}$. We know that $\mathcal{K}^{+}\in\mathbb{Y}$, so $\mathbb{Y}$ is non empty, hence there exists a containment minimal building set $B\in\mathbb{Y}$ such that $A\subsetneq B$. By the supersolvability of $\mathbb{B}(\mathcal{K})$ applied to the pair of building sets $A$ and $B$, we know that for any linear extension << of $\mathcal{K}$ and $x=\text{min}_{<}(B\setminus A)$, we have $B\setminus\{x\}\in\mathbb{B}(\mathcal{K})$. By our minimality assumption on $B$, we have that $(B\setminus\{x\})\cap\mathcal{L}^{+}\notin\mathbb{B}(\mathcal{L})$, i.e. $x$ is not an extreme point of $B$ in $\mathcal{L}$. We know that $I(\mathcal{L})\subseteq A$, hence $x\notin I(\mathcal{L})$ and $I(\mathcal{L})\subseteq(B\setminus\{x\})\cap\mathcal{L}^{+}$. By Proposition 3.16, there must exist some $y,z\in(B\setminus\{x\})\cap\mathcal{L}^{+}$ such that $y\wedge z\neq\hat{0}$ and $y\vee^{\mathcal{L}}z=x$. We know that $B\setminus\{x\}\in\mathbb{B}(\mathcal{K})$. Take $\{w_{1},\ldots,w_{k}\}=\text{max}\,(B\setminus\{x\})_{\leq x}$, and recall that for each $1\leq i<j\leq k$, we have that $\hat{0},w_{i}]\cap[\hat{0},w_{j}]=\hat{0}$. Note that since $x\notin B\setminus\{x\}$, we must have $k\geq 2$. Also, if $u\in(B\setminus\{x\})_{\leq x}$ and $u\wedge w_{i}\neq\hat{0}$, then $u\leq w_{i}$. As $y,z\in B\setminus\{x\}$ and $\hat{0}\neq y\wedge z$, we know that $y\vee^{\mathcal{K}}z\in B\setminus\{x\}$. We may assume without loss of generality that $y\vee^{\mathcal{K}}z\leq w_{1}$. Let $\{x_{1},\ldots,x_{t}\}=\text{max}\,I(\mathcal{L})_{\leq x}\subseteq B\setminus\{x\}$. Suppose that for some $s$ with $1\leq s\leq t$, we have $x_{s}\wedge y=x_{s}\wedge z=\hat{0}$. Because $I(\mathcal{L})\in\mathbb{B}(\mathcal{L})$, we know that for $1\leq i\leq t$ there exist $y_{i},z_{i}\leq x_{i}$ such that $\vee_{1\leq i\leq t}^{\mathcal{L}}y_{i}=y$ and $\vee_{1\leq i\leq t}^{\mathcal{L}}z_{i}=z$. By assumption, $y_{s}=z_{s}=\hat{0}$, so $y\vee^{\mathcal{L}}z\leq\vee_{i\neq s}^{\mathcal{L}}x_{i}<\vee_{1\leq i\leq t}^{\mathcal{L}}x_{i}=x$, a contradiction. Thus for each $i$, the intersection $x_{i}\wedge y\neq\hat{0}$ or $x_{i}\wedge z\neq\hat{0}$, hence $x_{i}\wedge w_{1}\neq\hat{0}$. As $x_{i}\in(B\setminus\{x\})_{\leq x}$, we must have $x_{i}\leq w_{1}$, therefore $I(\mathcal{L})_{\leq x}\subseteq[\hat{0},w_{1}]$. Because $\mathcal{L}\hookrightarrow\mathcal{K}$ is consistent, $I(\mathcal{K})_{\leq x}\subseteq I(\mathcal{L})_{\leq x}$. We conclude that $I(\mathcal{K})_{\leq x}\subseteq[\hat{0},w_{1}]$, implying $k=1$, a contradiction.
@@
(2): Suppose for contradiction that $B\in\mathbb{B}(\mathcal{K})$, and $I(\mathcal{L})\subseteq B$, but $\mathcal{N}_{\mathcal{L}}(B\cap\mathcal{L}^{+})\nsubseteq\mathcal{N}_{\mathcal{K}}(B).$ Let $N\in\mathcal{N}_{\mathcal{L}}(B\cap\mathcal{L}^{+})$ such that $N\notin\mathcal{N}_{\mathcal{K}}(B)$. There exists some $x_{1},\ldots,x_{k}\in N$ which are incomparable, such that $y=\bigvee_{1\leq i\leq k}^{\mathcal{K}}x_{i}\in B$ but $x=\bigvee_{1\leq i\leq k}^{\mathcal{L}}x_{i}\in\mathcal{L}^{+}\setminus B$. Let $\{y_{1},\ldots,y_{l}\}=\text{max}\,B_{\leq x}$. Then $l\geq 2$ and the intervals $\hat{0},y_{i}]$ intersect at $\hat{0}$. By Lemma 2.4, $y\leq x$, thus there exists exactly one $i$ such that $y\leq y_{i}$. But then by Lemma 2.10 applied to $x_{1},\ldots,x_{k}$, we know that for any $x^{\prime}\in(B\cap\mathcal{L}^{+})_{\leq x}$, there exists some $x_{t}$ such that $x^{\prime}\leq x_{t}$, hence $x^{\prime}\leq y\leq y_{i}$. It follows that for any $y_{j}$ with $j\neq i$, we have that $(B\cap\mathcal{L}^{+})_{\leq y_{j}}=\emptyset$. On the other hand, there must exist some $y^{\prime}\in I(\mathcal{K})_{\leq y_{j}}$. Clearly, $I(\mathcal{K})_{\leq y_{j}}\subseteq I(\mathcal{K})_{\leq x}$, and by our assumption that $\mathcal{L}\hookrightarrow\mathcal{K}$ is consistent, we must have that $y^{\prime}\in I(\mathcal{L})_{\leq x}\subseteq(B\cap\mathcal{L}^{+})_{\leq y_{i}}$, a contradiction.
@@ 5. Acknowledgements
We thank Nate Bottman, Federico Castillo, Graham Denham, Alexei Oblomkov, Arnau Padrol, Vincent Pilaud, Dasha Poliakova, and Raman Sanyal for helpful discussions about building sets, and we thank Bryan Gillespie for carefully explaining his work during the 2017 MSRI Semester on “Geometric and Topological Combinatorics”. Some of our discussions took place at the Simons Center for Geometry and Physics Workshop on “Combinatorics and Geometry of Convex Polyhedra” in March 2023. S.B. thanks the MSRI (now SLMath) and the Simons Center for their hospitality and excellent working conditions. We thank Christin Bibby, Graham Denham, Matt Larson, Arnau Pardol, Vincent Pilaud, Vic Reiner, and Raman Sanyal for their helpful feedback on an earlier version of this article.